Four liters of a 25% cleaning solution is mixed with a 65% solution to achieve a 45% solution? How much of teh 65% solution is needed?
Answer by stanbon(75887)

Four liters of a 25% cleaning solution is mixed with a 65% solution to achieve a 45% solution? How much of teh 65% solution is needed?
---------------
One-variable solution:
Equation:
active + active = active
0.25*4 + 0.65x = 0.45(4+x)
Multiply thru by 100 to get:
25*4 + 65x = 45*4 + 45x
20x = 20*4
x = 4 liters (amt. of 65% solution needed in the mixture)
=====================
Cheers,
Stan H.